A program executes 10 billion instructions. It exe
Answer #1

Given that

                         CPI = 1.5

                         Frequency = 2 GHz = 2 x 109

clock time = \frac {1}{frequency} = \frac {1}{2 \times 10^{9}} = 0.5 \times 10^{-9} = 0.5 ns

execution time = CPI \times Instruction count \times clock time (in secs)

                                = 1.5 \times (1 \times 10^{10}) \times (0.5 \times 10^{-9})

                                = 1.5 \times 1 \times 0.5 \times 10

                                = 7.5 secs

Thus, the given program with 10 billion instructions would take 7.5 seconds to execute on a intel processor with a clock period of 0.5 ns.

speed up = \frac {execution \: time\: with \: AMD \: processor}{execution \: time \: with \: Intel \:processor}

                   = \frac {6}{7.5} = 0.8

                   Hence, the intel processor has shown a 0.8 times speed up relative to AMD processor.